Obtain the expression for ?c as labelled in the following circuit valid for all t>0.

100 ? 40 ?F 30? 90 mH 9 V
Answer #1

Before t=0, the circuit is under steady state. Under stead state the capacitor acts like open circuit and the inductor will act like short circuit.

The circuit looks lke in the figure below.

100 Ω 9 V 30 Ω 90 mH

The voltage across 30ohm resister is 9V. Current through the resister is obtained by using ohms law

V = iR\Rightarrow i = \frac{V}{R}=\frac{9}{30} = 0.3V

Since capacitor is acting like an open circuit there is no current in 100 ohms resistatnce, so there is no voltage drop across it. Under steady state condition the inductor will act like a short circuit hence the voltage drop across the inductor is zero.

Writing a KVL in the loop

9-v_c-2i-v_{100\Omega}-v_{L}=0\Rightarrow 9-v_c-2i=0

v_c=9-2i = 9-2\times 0.3 = 9-0.6=8.4V

Hence the voltage across the capacitor is 8.4V